40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Tue Aug 18 15:05:33 2020, aaron, Update, Lab Work, notice of lab entry 
    Reply  Tue Aug 18 21:35:49 2020, aaron, Update, Lab Work, notice of lab entry 
Message ID: 2555     Entry time: Tue Aug 18 21:35:49 2020     In reply to: 2554
Author: aaron 
Type: Update 
Category: Lab Work 
Subject: notice of lab entry 
I'm taking inventory of our optics supply, in between figuring out what's wrong with the temperature log channels.

The inventory is in Clickup, which is a new organizational tool I'm trying out. There' an easy csv export, so I can get it elsewhere if/when we want. We have a wide variety of lenses:

Focal length (mm) 25.4 38.1 100 103.22 154.82 200 225 250 258.04 309.65 412.86 500 850
Quantity @ 1550nm 1 4 2 7 2 2 6 2 1 3 12 1 2

I found fewer curved mirrors, but there were a couple.

Steps I took for the temperature sensor:

1. Tried to see what's the temperature by reading the current temperature in ndscope or dataviewer. In dataviewer, go to 'Signal' tab, and enter the channel name or find it on the list of slow channels. For ndscope...

nsdscope X1:AUX-LAB_TEMP_F 

No data appears on either. I restart cymac1, which seemed frozen, but still nothing.

2. What's going on with these channels?

$ caget X1:AUX-LAB_TEMP_F
X1:AUX-LAB_TEMP_F     -459.7
$ caget X1:AUX-LAB_TEMP_K
X1:AUX-LAB_TEMP_K     -0.0166677

Looks like the channel is reading zero.

3. I traced the cable from the particle counter and found that it sends data to cominaux, the common auxiliary machine for the lab.

ssh cominauxThat's a little surprising to me, because
cd ~/services/modbusIOC/
less CRYOXT.db

This is the database file that defines the channels on cominaux. I search for 'LAB_TEMP_F' and find the epics record for the temperature channels. The epics records are all "calc" records, and the temperature in Kelvin is taken from X1:AUX-ACROXT_AI_15. This corresponds to channel 15 of the acromag slow ADC.

$ caget X1:AUX-ACROXT_AI_15
X1:AUX-ACROXT_AI_15     -0.0003333

 That's starting to make sense, the cable from the particle counter didn't go to the acromag ADC. Starting from the ADC channel 15, I traced the cable back to what used to be the AD590 temperature transducer.

4. Where did the IC temperature sensor go? Searching the elog and my dusty memory... neither readily recalls where it went. Let's get another one, they are cheap and easy to use.

ELOG V3.1.3-